OpenStudy (anonymous):

Translate the following algebraic expression: The sum of a number and two is three more than twice the same number x + 3 = 2 + 2x 2x + 2 = x + 2 3x + 3 = 2x + 2 x + 2 = 3 + 2x

OpenStudy (anonymous):

x + 2 = 3 + 2x

OpenStudy (anonymous):

x + 2 = 3 + 2x